OpenStudy (anonymous):

Lecture #10 homework check. Most of the textbook questions do not have answers in the back of the book, so I want to check my answers against what you got to see if I messed up.

OpenStudy (anonymous):

\[\text{A&E 9-76) } C^+\space \underset{s}{\fbox{↑↓}}\space\underset{p}{\fbox{↑ }\fbox{ }\fbox{ }}\space\underrightarrow{promotion}\space \underset{sp^2}{\fbox{↑ }\fbox{↑ }\fbox{↑ }}\space\underset{p}{\fbox{ }}\space\underrightarrow{bonding}\space \underset{sp^2}{\fbox{↑↓}\fbox{↑↓}\fbox{↑↓}}\space\underset{p}{\fbox{ }}\space CH^+_3\]Fancy markup, but I'm not sure about the promotion sequence on this one. I think the polyatomic ion is trigonal planar, which is different from methane. A&E 9-82) The square pyramidal sulfur pentafluoride is isoelectric with sulfur hexafluoride and has the following hybridization:\[sp^3d^2\] http://www.dur.ac.uk/a.k.hughes/vsepr/data/sf5_minus.html http://en.wikipedia.org/wiki/Sulfur_hexafluoride A&E 9-84) Diagram omitted. Bond order is 2. 1 unpaired electrons.\[[F=F]^{2+}\] A&E 9-88) Diagrams omitted. The types of orbitals and their tally is below.\[6\space MO \rightarrow \sigma+2\pi+2\pi^*+\sigma^*\] A&E 9-89) 2 antibonding electrons; X=X; bond distance increases A&E 9-95) Diagram omitted. Bond order 1. Stable.\[[O-O]^{2-}; \space [He]2s^22p^5\] A&E 9-96) Diagram omitted. Bond order 3. Stable.\[[C \equiv C]^{2-}; \space [He]2s^22p^3\] A&E 9-100) Diagram omitted. Bond order 2. Stable.\[B = \underset{\text{ . .}}{\overset{\text{ . .}}{N}} \quad \text{It does form. Is this what we are looking at?}\]